OpenStudy (anonymous):

It takes 15 minutes to walk to the mall from your house at an average velocity of 1.1 m/s to the east. What is the displacement from your house to the mall?

OpenStudy (anonymous):

Ok.. let's see.. We know that we walked for 15 minutes, avg speed 1.1m/s so.. \[ \text{length} = 1.1_{m/s} \cdot 15_{minutes} = 1.1_{m/s} \cdot (60\cdot 15)_{s} = 1.1_{m/s} \cdot 90_s = 99_m \]